OpenStudy (okdutchman7):

Please help. WILL MEDAL! @jim_thompson5910 @Kainui @sleepyjess @Ashleyisakitty @zepdrix @robtobey

OpenStudy (okdutchman7):

If \[f(x) = x ^{2} -x\] find \[\frac{ f(x+h) -f(x-h) }{ 2h }\]

OpenStudy (okdutchman7):

\[\frac{ (x ^{2}+h ^{2}+hx-(x+h))-(x ^{2}+h ^{2}-hx-9x-h)) }{ 2h }\]

jimthompson5910 (jim_thompson5910):

(x+h)^2 does not turn into x^2+h^2

OpenStudy (okdutchman7):

May you please explain?

jimthompson5910 (jim_thompson5910):

(x+h)^2 = (x+h)*(x+h) use FOIL or the box method